# DIsoN: Decentralized Isolation Networks for Out-of-Distribution Detection in Medical Imaging

### v1

@@ 4.1 Evaluation on Medical OOD Benchmarks
Tab. 1 compares CC-DIsoN with the baselines on our three medical OOD benchmarks where the OOD task is to detect artifacts. We can see that the post-hoc methods struggle to detect these domain-specific artifacts: fDBD and MDS average only $\sim 63\%$ AUROC while having a high FPR95 ($90.2\%$ and $74.6\%$, respectively). Training-time regularization methods like CIDER do better ($74.0\%$ average AUROC), but still have a high FPR95 of $71.6\%$. CC-DIsoN performs strongly compared to the baselines and shows consistent improvement across all datasets. Compared to MDS, it improves AUROC by 16.5% and reduces FPR95 by 18.4%. Against the best baseline (CIDER), it improves AUROC by 5.5% and lowers FPR95 by 15.4%. This demonstrates that directly comparing test samples against the training data during inference has a positive impact on OOD detection.
Table 3: Comparison of DIsoN and CC-DIsoN for three medical datasets.
AUROC $\uparrow$ FPR95 $\downarrow$
Dataset DIsoN CC-DIsoN DIsoN CC-DIsoN
Ultrasound 67.0 66.6 78.6 67.9
Dermatology 86.4 88.0 50.0 40.1
X-Ray 81.4 83.9 73.6 60.5
Average 78.3 79.5 67.4 56.2
@@
Effects of Class-Conditioning. Tab. 3 shows the benefits of class-conditioning (Sec. 3.3). On average, CC-DIsoN improves AUROC by 1.2%. More notably, CC-DIsoN reduces FPR95 by 11.2% across all three datasets. Class-conditioning consistently lowers FPR95, demonstrating that focusing the isolation task on the predicted same-class samples improves OOD detection.
Histopathology: MIDOG. While the earlier benchmarks focused on detecting non-diagnostic artifacts as OOD task, MIDOG evaluates semantic and covariate shift as OOD task. Tab. 2 shows that CC-DIsoN also performs strongly in this setting. In the most challenging near-OOD setting, CC-DIsoN achieves an average AUROC of 69.0%, which is a 4.5% improvement compared to the best regularization-based method PALM and a 4.9% gain compared to the best post-hoc method ViM. It also achieves the lowest FPR95 in this setting, albeit with a smaller margin. In the less challenging far-OOD setting, CC-DIsoN achieves an AUROC of 98.3% and a FPR95 of 5%, indicating almost perfect separation. Out of all compared methods, only PALM achieves better performance in the far-OOD task, although CC-DIsoN performs better in the near-OOD task. Overall, these results demonstrate that CC-DIsoN does not only perform well for OOD tasks with localised artifacts but also effective in identifying semantic and covariate shifts across medical imaging tasks.
@@
4.2 Further Analysis of DIsoN
Sensitivity Analysis of Hyperparameter $\alpha$. We analyze the effect of the aggregation weight $\alpha$ (Eq. 2) on both the OOD detection performance and convergence rate across Dermatology, Ultrasound, and X-Ray in Fig. 4(a). We can see that lower $\alpha$ values, which give relatively more weight to the target updates, reduce OOD detection performance, since the isolation task becomes dominated by the target sample, and loses the comparison signal to the source data. Increasing $\alpha$ improves AUROC consistently across datasets, with performance plateauing at $\alpha=0.8$. This shows that a stronger emphasis on training data improves the isolation-based OOD performance. However, the number of communication rounds required for convergence increases with $\alpha$. This aligns with Proposition 3.1, where $\alpha$ controls the implicit oversampling ratio N: smaller $\alpha$ increases the target signal (larger N) (faster isolation), while larger $\alpha$ slows convergence (more rounds needed to incorporate the target’s signal). In practice, this creates a trade-off between convergence speed and detection performance.
Effect of Network Size. To quantify how capacity of the Isolation Network affects the OOD scores, we compare three network sizes: a “Slim” ResNet18 (0.5× channel widths), the standard ResNet18, and a deeper ResNet34. Fig. 4(b) shows each backbone’s AUROC vs. FPR95 trade-off on Breast Ultrasound (circles), Dermatology (squares), and Chest X-Ray (triangles). Neither the Slim-ResNet18 nor the ResNet34 outperforms the standard ResNet18: the slim model might lack sufficient representational capacity to detect the subtle visual difference, while the deeper network does not yield consistent OOD gains and roughly doubling the required compute. This is likely because a binary isolation task does not require this excessive parameter capacity. Therefore, ResNet18 gives the best balance between good ID/OOD separation and computational efficiency.
Effect of Image Augmentations. Fig. 5 shows the effects of applying image augmentations during DIsoN training for Ultrasound, Dermatology, and X-Ray. Applying augmentations consistently improves AUROC: +4.43% on Ultrasound, +13.69% on Dermatology, and +2.76% on X-ray. FPR95 also decreases, most notably on Dermatology -30.99%. The density plot of dermatology dataset shows that without augmentation, ID and OOD samples converge in similar number of few rounds, resulting in poor separation. With augmentations, OOD target samples isolate quickly, but ID target samples require many more rounds. This demonstrates the regularization effect of augmentations: they prevent the model from quickly memorizing a single target sample, regardless of whether it is ID or OOD, and promote learning invariant features.
@@ Acknowledgements
FW is supported by the EPSRC Centre for Doctoral Training in Health Data Science (EP/S02428X/1), the Anglo-Austrian Society, and Oxford-Reuben scholarship. PS is supported by the EPSRC Programme Grant [EP/T028572/1], UKRI grant [EP/X040186/1], and EPSRC Doctoral Training Partnership award. HA is supported by a scholarship via the EPSRC Doctoral Training Partnerships programme [EP/W524311/1, EP/T517811/1]. The authors acknowledge the use of the University of Oxford Advanced Research Computing (ARC) facility (http://dx.doi.org/10.5281/zenodo.22558).
@@ Appendix A DIsoN and CC-DIsoN: Algorithm
Algorithm 1 Source Node: DIsoN / CC-DIsoN
@@
Algorithm 2 Target Node: DIsoN / CC-DIsoN
@@
where $g_{S}(\theta)=\frac{1}{|B_{s}|}\sum_{\mathbf{x}_{s}\in B_{s}}\nabla_{\theta}L(\theta;\mathbf{x}_{s},0)$, $g_{T}(\theta)=\nabla_{\theta}L(\theta;\mathbf{x}_{t},1)$ and $N$ the number of times $\mathbf{x}_{t}$ is oversampled in the centralized version.
@@
[…]
@@
Chest X-Ray Cardiomegaly 1788 1711 77 pacemaker 1000 224×224
@@
Histopathology Mitotic 421 375 46 near-OOD: domains 2–7 far-OOD: CCAgT & FNAC 500 per domain 50×50
@@ Appendix D Training Details & Hyperparameters
Dermatology Nevus vs. Non-Nevus ResNet18 Adam 750 $1\times 10^{-3}$ 32
@@
Chest X-Ray Cardiomegaly vs. No Cardiomegaly ResNet18 Adam 500 $1\times 10^{-3}$ 32
@@
Chest X-Ray ResNet18 Adam $3\times 10^{-3}$ 16
@@ D.2 Training DIsoN
As described in Section 3, we apply standard stochastic image augmentations across all dataset. The augmentations for each dataset are mostly identical, with only minor dataset-specific adjustments. For Histopathology, due to the smaller image size, we reduce the rotation range and leave out random cropping. For Breast Ultrasound, we replace random cropping with color jitter. The full set of augmentations are listed below:
  • Random rotation: ±$\pm 15^{\circ}$ (use ±$\pm 5^{\circ}$ for Histopathology)
  • Random crop: $224\times 224$ with padding=25 (applied to Chest X-Ray & Dermatology)
@@
Compute Time.
A full experiment run takes approximately 32 GPU hours for Chest X-ray, 7 GPU hours for Dermatology, 1–2 GPU hours for Breast Ultrasound, and 8 GPU hours per domain for Histopathology.
Appendix E Ablation Study: Further Investigation of Image Augmentations
@@
In Section 4.2, we showed that applying image augmentations during DIsoN training improves OOD detection across Dermatology, Chest X-ray, and Breast Ultrasound. Here, we extend this ablation study by comparing four augmentation settings: no augmentation, augmentations only on the Source Node data, only on the target sample, and on both nodes. Fig. 6 reports AUROC and FPR95 for all settings. Applying augmentations on both nodes consistently performs best. Interestingly, for Dermatology, applying augmentations only on the target sample gives best AUROC overall. We can also see that target-only augmentation outperforms source-only augmentations in most cases. One exception is Ultrasound FPR95, where source-only yields better results. These findings further strengthen our finding of the importance of regularization on especially the target sample to prevent fast overfitting and to encourage learning of invariant features, as discussed in Section 4.2.
@@
Appendix F Qualitative Visualization of DIsoN Isolation Process

### v2

@@ 4.1 Evaluation on Medical OOD Benchmarks
Tab. 1 compares CC-DIsoN with the baselines on our three medical OOD benchmarks where the OOD task is to detect artifacts. We can see that the post-hoc methods struggle to detect these domain-specific artifacts: fDBD and ViM average only between $63-64\%$ AUROC while having a high FPR95 ($84.6\%$ and $78.1\%$, respectively). Training-time regularization methods like CIDER do better ($70.2\%$ average AUROC), but still have a high FPR95 of $72.8\%$. CC-DIsoN performs strongly compared to the baselines and shows consistent improvement across all datasets. Compared to fDBD, it improves AUROC by 15.8% and reduces FPR95 by 25.4%. Against the best baseline (CIDER), it improves AUROC by 9.8% and lowers FPR95 by 13.6%. This demonstrates that directly comparing test samples against the training data during inference has a positive impact on OOD detection.
@@
Effects of Class-Conditioning. Tab. 3a shows the benefits of class-conditioning (Sec. 3.3). On average, CC-DIsoN improves AUROC by 1.7%. More notably, CC-DIsoN reduces FPR95 by 8.2% across all three datasets. Class-conditioning consistently lowers FPR95, demonstrating that focusing the isolation task on the predicted same-class samples improves OOD detection.
Table 3: (a) Comparison of DIsoN and $C$-DIsoN on three medical datasets using AUROC (higher is better) and FPR95 (lower is better). (b) Effect of incorrect predicted classes for class-conditional sampling in $C$-DIsoN (AUROC). “Best Baseline” is the strongest baseline from Tab. 1 (for reference); “ID & OOD wrong”: all targets assigned incorrect classes; “ID wrong”: only ID targets assigned incorrect classes; “OOD wrong”: only OOD targets assigned incorrect classes.
Dataset AUROC $\uparrow$ FPR95 $\downarrow$
DIsoN $C$-DIsoN DIsoN $C$-DIsoN
Ultrasound 67.0 65.6 78.6 73.2
Dermatology 86.4 89.5 50.0 42.5
X-Ray 81.4 84.9 73.6 61.9
Average 78.3 80.0 67.4 59.2
Dataset Best Baseline Standard CC-DIsoN ID & OOD wrong ID wrong OOD wrong
Ultrasound 61.0 65.6 61.5 34.6 89.6
Dermatology 82.0 89.5 83.1 84.7 86.7
X-Ray 70.5 84.9 77.9 74.0 88.3
Impact of Predicted Class Accuracy. Since CC-DIsoN conditions on the predicted class from the pre-trained model $M_{pre}$, we analyse how misclassification affect its performance. We simulate three controlled scenarios by asssigning wrong predicted classes to: (i) all target samples (ID & OOD), (ii) only ID targets while OOD targets use their predicted classes, and (iii) only OOD targets while ID targets use their predicted classes. Results in Tab.3b show that performance drops mainly when ID targets are incorrect, as class-conditioning compares them to unrelated classes, making isolation easier and reducing the separation in convergence rounds between ID and OOD samples. In contrast, performance slightly increases when OOD targets are mislabeled, as comparing them to unrelated classes makes convergence faster. Even under such extreme settings, CC-DIsoN remains competitive. Further details are provided in the Appendix E.
@@
Histopathology: MIDOG. While earlier benchmarks focused on detecting non-diagnostic artifacts as OOD task, MIDOG evaluates semantic and covariate shift as OOD task. Tab. 2 shows that CC-DIsoN also performs strongly in this setting. In the most challenging near-OOD setting, CC-DIsoN achieves an average AUROC of 69.6%, which is a 8.4% improvement compared to the best regularization-based method PALM and a 7.2% gain compared to the best post-hoc method ViM. It also achieves the lowest FPR95 in this setting, albeit with a smaller margin. In the less challenging far-OOD setting, CC-DIsoN achieves an AUROC of 98.3% and a FPR95 of 4.4%, indicating almost perfect separation. Out of all compared methods, only PALM achieves same AUROC performance in the far-OOD task, although CC-DIsoN performs better in the near-OOD task. IForests were originally developed for tabular data, where they perform well. However, prior works  [34, 10] shows poor performance on high-dimensional imaging data, which explains the low AUROC in our experiments. Overall, these results demonstrate that CC-DIsoN does not only perform well for OOD tasks with localised artifacts but also effective in identifying semantic and covariate shifts across medical imaging tasks.
4.2 Further Analysis of DIsoN
Sensitivity Analysis of Hyperparameter $\alpha$. We analyze the effect of the aggregation weight $\alpha$ (Eq. 2) on both the OOD detection performance and convergence rate across Dermatology, Ultrasound, and X-Ray in Fig. 4(a). We can see that lower $\alpha$ values, which give relatively more weight to the target updates, reduce OOD detection performance, since the isolation task becomes dominated by the target sample, and loses the comparison signal to the source data. Increasing $\alpha$ improves AUROC consistently across datasets, with performance plateauing at $\alpha=0.8$. This shows that a stronger emphasis on training data improves the isolation-based OOD performance. However, the number of communication rounds required for convergence increases with $\alpha$. This aligns with Proposition 3.1, where $\alpha$ controls the implicit oversampling ratio N: smaller $\alpha$ increases the target signal (larger N) (faster isolation), while larger $\alpha$ slows convergence (more rounds needed to incorporate the target’s signal). In practice, this creates a trade-off between convergence speed and detection performance. The goal of this analysis is not to identify one “optimal” value of $\alpha$, but to show that our method remains robust across a wide range of values. Performance stays quiet stable for $\alpha$ between 0.5 and 0.95 (Fig. 4a), suggesting good generalization across tasks without extensive hyperparameter tuning.
@@ 4.2 Further Analysis of DIsoN
Effect of Network Size. To quantify how capacity of the Isolation Network affects the OOD scores, we compare three network sizes: a “Slim” ResNet18 (0.5× channel widths), the standard ResNet18, and a deeper ResNet34. Fig. 4(b) shows each backbone’s AUROC vs. FPR95 trade‐off on Breast Ultrasound (circles), Dermatology (squares), and Chest X-Ray (triangles). Neither the Slim‐ResNet18 nor the ResNet34 outperforms the standard ResNet18: the slim model might lack sufficient representational capacity to detect the subtle visual difference, while the deeper network does not yield consistent OOD gains and roughly doubling the required compute. This is likely because a binary isolation task does not require this excessive parameter capacity. Therefore, ResNet18 gives the best balance between good ID/OOD separation and computational efficiency.
Effect of Image Augmentations. Fig. 5 shows the effects of applying image augmentations during DIsoN training on Ultrasound, Dermatology, and X-Ray. Augmentations consistently improves AUROC: +4.43% on Ultrasound, +13.69% on Dermatology, and +2.76% on X-ray. FPR95 also decreases, most notably on Dermatology -30.99%. The density plot of dermatology dataset shows that without augmentation, ID and OOD samples converge in similar number of few rounds, resulting in poor separation. With augmentations, OOD target samples isolate quickly, but ID target samples require many more rounds. This demonstrates the regularization effect of augmentations: they prevent the model from quickly memorizing a single target sample, regardless of whether it is ID or OOD.
Runtime Analysis and Practical Considerations. We extend the runtime analysis in Fig. 4a by reporting detailed statistics, including quantiles and results across different $\alpha$ values, in Appendix G. Since wall-clock time depends on hardware and network conditions, we primarily measure runtime in communication rounds, offering a hardware-independent estimate. Although DIsoN introduces extra computation, as each target sample requires an isolation task, this design trades speed for improved OOD detection performance. In practice, real-time inference is rarely required in healthcare, where scans are often reviewed hours or days later [31] (except in emergencies), and diagnostic accuracy is typically prioritised over speed. Further runtime comparisons with baselines, as well as notes on parallelisation and more detailed practical considerations, are provided in Appendix H.
@@ Acknowledgements
FW is supported by the EPSRC Centre for Doctoral Training in Health Data Science (EP/S02428X/1), the Anglo-Austrian Society, and an Oxford-Reuben scholarship. PS is supported by the EPSRC Programme Grant [EP/T028572/1], UKRI grant [EP/X040186/1], and EPSRC Doctoral Training Partnership award. HA is supported by a scholarship via the EPSRC Doctoral Training Partnerships programme [EP/W524311/1, EP/T517811/1]. The authors acknowledge the use of the University of Oxford Advanced Research Computing (ARC) facility (http://dx.doi.org/10.5281/zenodo.22558).
@@ Appendix A DIsoN and CC-DIsoN: Algorithm
Algorithm 1 Source Node: DIsoN / $C$-DIsoN
@@
Algorithm 2 Target Node: DIsoN / $C$-DIsoN
@@
where $g_{S}(\theta)=\frac{1}{|B_{s}|}\sum_{\mathbf{x}_{s}\in B_{s}}\nabla_{\theta}L(\theta;\mathbf{x}_{s},0)$, $g_{T}(\theta)=\nabla_{\theta}L(\theta;\mathbf{x}_{t},1)$ and $N$ is the number of times $\mathbf{x}_{t}$ is oversampled in the centralized version.
@@
$=\alpha\bigl(\theta^{(r)}-\eta\,g_{S}\bigr)+\beta\bigl(\theta^{(r)}-\eta\,g_{T}\bigr)\quad\text{(by (A.2))}$
@@
Chest X‐Ray Cardiomegaly 1788 1711 77 pacemaker 1000 224×224
@@
Histopathology Mitotic 421 375 46
near‐OOD: domains 2–7
far‐OOD: CCAgT & FNAC
500
per domain
50×50
@@ Appendix D Training Details & Hyperparameters
Dermatology Nevus vs. Non‐Nevus ResNet18 Adam 750 $1\times 10^{-3}$ 32
@@
Chest X‐Ray Cardiomegaly vs. No Cardiomegaly ResNet18 Adam 500 $1\times 10^{-3}$ 32
@@
Chest X‐Ray ResNet18 Adam $3\times 10^{-3}$ 16
@@ D.2 Training DIsoN
As described in Section 3, we apply standard stochastic image augmentations across all datasets. The augmentations for each dataset are mostly identical, with only minor dataset-specific adjustments. For Histopathology, due to the smaller image size, we reduce the rotation range and leave out random cropping. For Breast Ultrasound, we replace random cropping with color jitter. The full set of augmentations are listed below:
  • Random rotation: $\pm 15^{\circ}$ (use $\pm 5^{\circ}$ for Histopathology)
  • Random crop: $224\times 224$ with padding=25 (applied to Chest X‐Ray & Dermatology)
@@
Convergence Parameter Choices for $E_{\mathrm{stab}}$ and $\tau$.
The parameters $E_{\mathrm{stab}}$ and $\tau$ were chosen to capture the convergence behavior of the isolation process. Specifically, $E_{\mathrm{stab}}$ acts as a patience parameter, conceptually similar to early-stopping criteria in learning-rate schedulers, requiring correct classification of $\mathbf{x}_{t}$ for several consecutive rounds to ensure stable convergence. We set $E_{\mathrm{stab}}=5$, a commonly used value in schedulers that effectively captures convergence. The confidence threshold $\tau=0.85$ ensures that the model learned to separate the target sample with sufficient confidence. This value follows standard practice in uncertainty-based literature [18]. In preliminary experiments, these values consistently produced stable and reliable results, so they were kept fixed across all datasets to avoid dataset-specific tuning or overfitting.
Effect of partial Fine-Tuning
We also examined whether freezing parts of the network could be beneficial for the isolation task by simplifying optimization. Freezing the backbone, however, assumes that the pre-trained feature extractor already provides sufficiently expressive embeddings to distinguish ID from OOD samples. We found this assumption too restrictive: in early experiments, partial fine-tuning of the network (e.g., only the head or last block) led to worse OOD detection performance than fine-tuning the entire model.
Appendix E Additional Details: Ablation Study for Impact of Predicted Class Accuracies for Class-Conditional Sampling
CC-DIsoN uses predictions of a primary model $M_{\text{pre}}$, deployed for a task of interest (e.g., cardiomegaly classification in X-rays; see Sec. 4), for class-conditional sampling. This enables comparing a target sample to the most relevant subset of training data (its predicted class) rather than to all training samples. Our ablations on the effects of class-conditioning (Tab. 3a) demonstrated that class-conditional sampling improves OOD detection performance, while DIsoN without class-conditioning still performs well (see comparison between Tabs. 3a and 1).
To further analyze the influence of class-conditioning and how misclassifications affect the results, we provide here a more detailed analysis of the ablation study on predicted class accuracy introduced in Section 4.1. We first measure the classification accuracy of $M_{\text{pre}}$ on ID, OOD, and combined target samples, along with the corresponding CC-DIsoN AUROC scores (Tab. 7). Typically, especially in safety-critical domains such as healthcare, high ID accuracy is a requirement for deployment. It is also expected that classification accuracy decreases on OOD samples due to domain shift. Comparing these accuracies with OOD detection performance shows no clear correlation, for example, the Dermatology dataset achieves the highest AUROC (89.5) despite the lowest classification accuracy (63.0%). This indicates that CC-DIsoN’s OOD detection performance does not depend on perfectly accurate predicted classes but rather on the data characteristics.
Table 7: Classification accuracy of the primary models on ID, OOD, and combined target samples, together with $C$-DIsoN OOD detection performance (AUROC from Tab. 1).
Dataset ID Acc. (%) OOD Acc. (%) ID&OOD Acc. (%) CC-DIsoN AUROC
Dermatology 77.6 54.0 63.0 89.5
Chest X-Ray 93.0 70.8 82.2 84.9
Ultrasound 78.6 61.4 64.8 65.6
MIDOG (near+far) 77.3 68.6 69.2 76.0
In Section 4.1 and Tab. 3b, we further investigate the effect of incorrect predicted classes on class-conditioning in controlled settings. Here, we describe these settings in more detail. We designed three experiment variants for Dermatology, Breast Ultrasound, and Chest X-ray by manually assigning wrong predicted classes from the primary model to specific target samples:
  1. 1.
    All targets mislabeled (ID + OOD wrong): All target samples, regardless of whether they are ID or OOD, are assigned a random incorrect class. This represents the extreme case of a completely inaccurate primary model (0% Accuracy), which would not be realistic in deployment but an interesting setting to obtain insights.
  2. 2.
    Only ID targets mislabeled: All target ID samples are assigned incorrect classes (0% Accuracy on ID), while OOD targets use their predicted classes. Again, this represents an impractical scenario for deployment, since a model with 0% ID accuracy would not be used in real-world, but useful to study misclassification behavior. Central motivation for class-conditioned sampling was to make the isolation task more difficult for ID target samples, as they should be harder to isolate from source (training ID) samples of the same class, as they share common visual features. This should lead to slower convergence and increase the difference than when separating OOD samples from source (training ID) samples, even of the same class, which is faster due to domain-shift. If we instead compare ID target samples to source (training) samples from the wrong class, the isolation will be easier, resulting in worse OOD detection.
  3. 3.
    Only OOD targets mislabeled: All OOD target samples are assigned incorrect classes (0% Accuracy on OOD), while ID targets use their predicted classes. This mimics a situation where domain shift causes the primary model to misclassify OOD samples, and gives us insights what would happen in the most extreme cases of domain shift. Since OOD data already differ visually from the source distribution, assigning and comparing them with unrelated classes can make their isolation even easier, and we expect an improvment in detection.
The results in Tab. 3b confirm these expectations. Performance decreases when ID targets are mislabeled, where class-conditioning is most important, while slightly increasing when OOD targets are mislabeled. Even in the extreme “all wrong” case, CC-DIsoN remains competitive and outperforms the strongest baseline (Tab. 3b) in all but one dataset. This shows that CC-DIsoN’s improvement does not rely on perfectly accurate class predictions and that the method remains robust to moderate number of misclassification errors. Overall, comparison with visually similar source samples is beneficial for DIsoN and provides stable OOD detection performance even when predicted classes are not perfect.
Appendix F Ablation Study: Further Investigation of Image Augmentations
@@
In Section 4.2, we showed that applying image augmentations during DIsoN training improves OOD detection across Dermatology, Chest X-ray, and Breast Ultrasound. Here, we extend this ablation study by comparing four augmentation settings: no augmentation, augmentations only on the Source Node data, only on the target sample, and on both nodes. Fig. 6 reports AUROC and FPR95 for all settings. Applying augmentations on both nodes consistently performs best. Interestingly, for Dermatology, applying augmentations only on the target sample gives best AUROC overall. We can also see that target-only augmentation outperforms source-only augmentations in most cases. One exception is Ultrasound FPR95, where source-only yields better results. These findings further strengthen our finding of the importance of regularization on especially the target sample to prevent fast memorization of superficial image-specific features, and encourage learning more meaningful features for differentiation of ID and OOD data, as discussed in Section 4.2.
@@
Appendix G Runtime Statistics
Figure 4 (a) in the main paper reports the average number of rounds required for the isolation task to converge. To provide a more detailed view of runtime variability, we report additional statistics in Table 8, including the 25th percentile, median, and 75th percentile of convergence rounds for both ID and OOD target samples using our default $\alpha=0.8$. Reporting these quantiles helps practitioners assess expected runtime under different ID/OOD ratios. Since wall-clock runtime depends on hardware, we report the number of rounds as a hardware-independent measure. OOD samples generally converge in fewer rounds than ID samples, consistent with their easier separability from the source distribution.
Table 8: Quantiles (25th, median, 75th percentile) of convergence rounds for ID and OOD samples across datasets. OOD samples converge in fewer rounds and show lower variability, while ID samples display higher variability due to within-distribution complexity.
Dataset Sample Type 25th Perc. Median 75th Perc.
Dermatology ID 68.0 126.5 184.0
OOD 22.0 28.0 40.0
Ultrasound ID 20.0 29.0 100.0
OOD 16.0 20.5 32.0
Chest X-Ray ID 28.0 41.0 92.0
OOD 16.0 20.0 26.0
Histopathology (near OOD) ID 44.3 59.5 74.0
OOD 34.0 44.0 57.0
We further analyze how the runtime distribution changes with $\alpha$. Tables 9-11 show that decreasing $\alpha$ consistently reduces the number of rounds required for convergence. Importantly, OOD samples converge faster across all $\alpha$, while ID samples exhibit broader variability due to within-distribution complexity. Note that, as described above, we set the maximum number of rounds $R_{\text{max}}$ as the OOD score when convergence was not reached. For these additional runtime experiments, we set a lower maximum number of rounds for the Dermatology dataset from ($R_{\text{max}}=300$) to ($R_{\text{max}}=150$) in order to reduce computational cost. For $\alpha=0.90$ and $\alpha=0.95$, the majority of ID samples did not converge before reaching this limit. Therefore, the 25th percentile equals $R_{\text{max}}$ in those cases.
Table 9: Distribution of convergence rounds (25th, median, 75th percentile) for Breast Ultrasound across $\alpha$ values. Lower $\alpha$ accelerates convergence, while OOD samples consistently require fewer rounds than ID samples.
$\alpha$ Sample Type 25th Perc. Median 75th Perc.
0.95 ID 39.8 50.0 300.0
OOD 33.0 42.0 59.0
0.90 ID 25.0 39.5 297.0
OOD 22.0 27.0 42.0
0.50 ID 13.8 17.0 26.3
OOD 13.0 15.0 19.0
0.40 ID 11.0 13.0 17.0
OOD 11.0 14.0 17.0
Table 10: Distribution of convergence rounds (25th, median, 75th percentile) for Chest X-Ray across $\alpha$ values. Lower $\alpha$ reduces the number of rounds required for convergence, with OOD samples converging faster and more consistently than ID samples.
$\alpha$ Sample Type 25th Perc. Median 75th Perc.
0.95 ID 51.0 58.0 75.0
OOD 44.0 48.0 53.0
0.90 ID 26.0 35.0 100.0
OOD 19.0 23.0 27.0
0.50 ID 20.0 27.0 51.5
OOD 14.0 16.0 20.0
0.40 ID 15.0 17.0 23.0
OOD 11.0 13.0 15.0
Table 11: Distribution of convergence rounds (25th, median, 75th percentile) for Dermatology across $\alpha$ values. Convergence becomes substantially faster as $\alpha$ decreases, and OOD samples consistently require fewer rounds than ID samples.
$\alpha$ Sample Type 25th Perc. Median 75th Perc.
0.95 ID 150.0 150.0 150.0
OOD 72.0 87.0 122.0
0.90 ID 150.0 150.0 150.0
OOD 37.0 44.0 66.0
0.50 ID 16.0 17.0 21.0
OOD 13.0 15.0 16.0
0.40 ID 5.0 11.0 15.0
OOD 5.0 5.0 11.0
Appendix H Practical Deployment Considerations And Runtime Comparisons
In many medical imaging workflows, real-time inference is not required. Diagnostic scans are often reviewed hours or days after acquisition due to limited clinical staff availability, and in some cases, turnaround times can extend to several weeks [31]. Within such workflows, AI models can operate asynchronously, processing scans during this waiting period. Therefore, inference times of several minutes per sample are acceptable when they provide more reliable detection of OOD samples, and similar latency is acceptable in many other non-real-time workflows outside of healthcare. DIsoN was specifically developed for such workflows, where reliability and privacy are prioritised over speed, and these latencies remain practically acceptable. To provide a complete view of computational cost, we report both communication rounds (Appendix G) and wall-clock runtimes for comparison with other methods. Together, these results help practitioners assess whether DIsoN is suitable for their deployment setting.
DIsoN introduces additional computation at inference time (approximately 40 s–4 min per sample in our experiments on an NVIDIA RTX A5000), as each target sample requires iterative isolation training until convergence. In contrast, most compared baselines complete inference within milliseconds, since they rely on a single forward pass of a CNN backbone followed by lightweight post-processing to compute an OOD score (e.g., distance computation, linear projection, or tree traversal). Note that iForest, conceptually the closest to DIsoN, cannot be trained jointly on source and target data due to privacy constraints, unlike DIsoN. Following the original iForest formulation [25], we pre-train it only on the source (ID) data and then apply it with a single forward pass per target sample to determine whether it is ID or OOD, which explains its much lower runtime.
While this makes the baselines faster, DIsoN’s iterative optimisation results in more reliable separation between ID and OOD samples, as demonstrated in our results. Average per-sample runtimes for representative baselines on the Ultrasound dataset are reported in Table 12.
Table 12: Average inference time per target sample (seconds, NVIDIA RTX A5000). Baseline methods compute OOD scores after a single forward pass of the pre-trained network followed by lightweight post-processing, whereas CC-DIsoN performs iterative optimisation until convergence, which increases latency but yields more reliable OOD detection performance.
Method Avg. Runtime (s)
MDS 0.006
ViM 0.005
iForest 0.015
CIDER 0.007
PALM 0.007
CC-DIsoN (ID samples) 155.2
CC-DIsoN (OOD samples) 90.9
Single-Sample vs. Batch Processing. DIsoN is designed to process one target sample at a time, as it learns a decision boundary that separates a single target sample from the training distribution using convergence behaviour as the OOD score. This design aligns with many real-world scenarios, especially in healthcare, where each patient scan is acquired and analysed individually rather than in large batches. If batch processing is feasible, multiple DIsoN instances can be executed in parallel to use available hardware efficiently. In our experiments, up to eight DIsoN runs could be executed concurrently on a single NVIDIA RTX A5000 (24 GB VRAM). Although not a traditional deep learning batch-tensor setup, this parallel execution allows scaling for pre-collected datasets.
Network Conditions and Communication Latency. DIsoN involves a single target client communicating with a source node, making it simpler and more robust than large-scale multi-client federated systems. Network delays or volatility only affect wall-clock time, not the number of communication rounds or the quality of convergence. In practice, network disconnections can be handled through standard retry or timeout mechanisms used in distributed systems.
Overall, these experiments and considerations demonstrate that while DIsoN trades inference speed for reliability, its runtime remains predictable and manageable across deployment conditions, making it suitable for safety-critical settings where performance is prioritised and real-time inference is not required.
Appendix I Additional Multi-Seed Results for Histopathology
Due to space limitations, Table 2 reports means over three seeds only, the standard deviations are omitted. In this section, we provide the complete tables including standard deviations. Tables 13 (AUROC $\uparrow$) and 14 (FPR95 $\downarrow$) list the mean $\pm$ std over three random seeds for the Histopathology dataset. For the Avg. columns, the $\pm$ values are computed as the sample standard deviation across seeds of the per-seed macro averages.
Table 13: Complete AUROC ($\uparrow$) results for the Histopathology dataset (MIDOG). Each cell shows mean $\pm$ standard deviation over three random seeds. This extends the means reported in Table 2 with standard deviation.
Methods near-OOD far-OOD
2 3 4 5 6a 6b 7 Avg. CCAgT FNAC Avg.
[…]
Table 14: Complete FPR95 ($\downarrow$) results for the Histopathology dataset (MIDOG). Each cell shows mean $\pm$ standard deviation over three random seeds. This extends the means reported in Table 2 with standard deviation.
Methods near-OOD far-OOD
2 3 4 5 6a 6b 7 Avg. CCAgT FNAC Avg.
[…]
Appendix J Qualitative Visualization of DIsoN Isolation Process